HOME ESR EXPERIMENTS LABS etc
E121 E127 E132 E125 E143 laser_cooling_2021 E142 Ê128 E146 E0052 E0018 E0028 E0038 E0075 E0203_E0018
  masses & isomeres 2025  Not logged in ELOG logo
Message ID: 52     Entry time: Tue Apr 1 22:03:07 2025
Author: Jeroen, Zac, Esther, Shahab, Wolfram 
Category: DAQ 
Subject: Optimisation of settings 
Estimation of potential gain by changing. Files from this points are going to rubbish. Potential gains are c. 20% 

22:04: The current acquisition window is recording for 2.9 s. And is not recording for 3.9 s. We are running currently with a spill cycle of 5s every repetition.

22:22 The spill cycle between each injection was changed to 4->3->2s progressively and FRS also stopped taking spills so we settled on 2s spill request.

In the morning the spill request has to be changed back to 5s when the FRS goes back into operation.

new file with the setting is

RSA01-2025.04.01.22.19.06.761.tiq

[analysis]
nframes = 190
lframes = 65536
navg = 1
ELOG V3.1.5-fc6679b